OpenStudy (anonymous):

(x + 1)(x - 1) i got x^2 - 2?

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

\[1\times -1\] is the last term

OpenStudy (anonymous):

what on earth?

OpenStudy (anonymous):

sorry but this is not right.

OpenStudy (anonymous):

no i know its not right

OpenStudy (anonymous):

imran will explain why. but the last term is not -2

OpenStudy (anonymous):

well even that is wrong. dang

OpenStudy (anonymous):

x*x-1*x+1*x -1*1

OpenStudy (anonymous):

x*x -1*x +1*x -1*1

OpenStudy (anonymous):

\[(x+1)(x-1)=x(x-1)+1(x-1)=x^2-x+x-1=x^2-1\]

OpenStudy (anonymous):

for example if \[x=10\] you know that \[(10+1)(10-1)=11\times 9=99=10^2-1\]